OpenStudy (anonymous):

Which method would you choose to solve this equation? Justify your reasoning. x^2+6X-2=0 I know that I want to graph this but I am not sure how to justify my reasoning.

OpenStudy (anonymous):

Complete the square?

OpenStudy (anonymous):

i would complete the square as well because the "middle term" 6x has an even coefficient

OpenStudy (anonymous):

then you do not get an annoying denominator forced on you by the quadratic formula

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

\[(x^2+6x=2\] \[(x+3)^3=2+9=11\] etc

OpenStudy (anonymous):

That should be (x+3)^2 though.

OpenStudy (anonymous):

yeah right! don't know where the cube came from...

OpenStudy (anonymous):

Is "completing the square" this method : x^2+6X-2=0 x² + 2 *3 x -2 =0 x² + 2 *3 x + 3² - 3² -2 =0 (x+3)² -11 = 0 (x+3)² - Sqrt(11)² = 0 (x+2 - Sqrt(11)) * (x+2+Sqrt(11)) = 0 x = -2 + Sqrt(11), x = -2 - Sqrt(11)

OpenStudy (anonymous):

Something like that, except I think you've overcomplicated it. From the fourth line you have:\[(x+3)^2=11\implies x+3=\pm\sqrt{11}\implies x=-3\pm\sqrt{11}\]

OpenStudy (anonymous):

Thanks .